# Navigating Heterogeneity and Privacy in One-Shot Federated Learning with Diffusion Models

### v1

@@ 4.1 Data Heterogeneity
In Table 1, we analyze the performance of all methods under moderate ($Dir(0.1)$) to extreme ($Dir(0.001$) heterogeneity. Interestingly, FedDiff outperforms all other methods by a significant margin, from $\sim$5to up to $\sim$20in different scenarios. In the case of CIFAR-10, which is the most complex of the datasets, we find that FedDiff provides the most improvement. As discussed in our initial motivations (Section 3), we reason that the focus on sample quality and diversity that is provided by the DM objective enables much improved performance. Intuitively, this becomes increasingly evident in more complex settings. To verify this observation, we conduct a comparative analysis of generated samples produced by FedCVAE and our FedDiff approach, illustrated in Figure 2. The discernible disparity is evident, with the samples generated by our method exhibiting significantly enhanced sharpness and overall quality.
(a) FedCVAE
(b) FedDiff (ours)
Figure 2: Random sets of generated samples from FedCVAE and our FedDiff approach. By leveraging the intrinsic properties of diffusion models (DMs), which are well-aligned with the requirements of one-shot FL, we achieve substantial benefits in sample quality and subsequent global model performance.
@@ 4.2 Number of Clients
Deepening our investigation, we also study the effect of the number of clients $C$ in Table 2. Note that, as we employ the same total number of samples in all experiments, the number of samples per client will increase with smaller $C$, and decrease with larger $C$. This allows us to observe the effect of increasing the distributed nature of the data across the client network.
One question arising from the adoption of generative models in FL settings pertains to their ability to maintain satisfactory performance when trained on a limited number of samples. Interestingly, when analyzing the results, we find that FedDiff is capable of handling a much smaller number of client training samples with little to no performance degradation. On the other hand, the discriminative model approaches quickly experience a collapse in performance when expanding to 20 clients. In the heterogeneous environment of federated learning, the local optimization of a discriminative model on a highly-imbalanced and small dataset proves challenging. Rather than being an overwelming burden, such a situation is handled well by FedDiff, as its sole focus is to capture the subsequently smaller distribution. Furthermore, we again find that FedDiff outperforms FedCVAE in all settings, further illustrating the potential for diffusion models in one-shot FL.
@@ 4.3 Resource Requirements
It is pertinent to emphasize that training diffusion models within the FedDiff framework is no more intricate than conventional methodologies and remains highly viable for FL. The computational complexity aligns with training a conventional CNN model with a modest number of parameters, and we employ the same number of local epochs as previous work with CNNs [35]. Importantly, the training process entails selecting random steps in the diffusion process at any given training iteration, eliminating the necessity for sequential steps during training. During inference, the generation process involves sequential denoising steps; however, this poses no issue for the clients, as generation occurs at the server in FedDiff. Therefore, FedDiff is an effective and practical approach for providing strong performance.
@@ 5 RQ2: Privacy Considerations
Privacy holds paramount importance in one-shot FL. The trained client model may be repeatedly utilized, or even exchanged in a model market context, and therefore safeguarding the privacy of the model before it leaves the client is imperative. However, other SOTA works have not experimented with DP constraints, nor have they thoroughly explored this aspect, often leaving privacy discussions simply as a possibility for future work [35, 13]. In the subsequent sections, we meticulously investigate privacy from various perspectives and delve into our research question RQ2.
@@ 5.1 Differential Privacy
Differential privacy is the widely accepted standard for ensuring privacy of a model, as it offers a provable guarantee of privacy [1, 8, 9, 7, 11]. Utilizing $(\epsilon,\delta)$ differential privacy during model training guarantees comprehensive privacy protection, encompassing not only the model’s parameters and activations, but also extending to all subsequent downstream operations such as inferences, fine-tuning, and distillation. It is important to note that a model trained under $(\epsilon,\delta)$ DP safeguards the privacy of every training sample, regardless of its qualities or uniqueness [34]. Specifically, we train all approaches under ($\epsilon,\delta$) DP at the clients for various privacy levels of $\epsilon$ = 50, 25, and 10, with $\delta=10e^{-5}$, $C=10$, and $\alpha=0.01$. Lower $\epsilon$ values correspond to a tighter privacy budget, and the stated budget is for the entire training of each local model. We employ the Opacus [34] library for implementing DP. Further DP training details are provided in Supplementary Material, along with additional $\epsilon$ experiment. We present the results for all approaches in Table 4.
@@
As expected, all methods experience a drop in performance when trained under DP settings. Nonetheless, FedDiff still stands out, outperforming all other methods by a significant margin. Particularly for FashionMNIST, FedDiff experiences comparatively less accuracy drop under DP than FedCVAE. As articulated in our initial motivations outlined in Section 3, DP training introduces noise into the training process, exacerbating the complexity of optimization. In such scenarios, the simplicity of the training paradigm employed by diffusion models becomes notably advantageous. Overall, we show that FedDiff is a strong approach even when DP is employed.
5.2 Addressing Memorization
In a recent study, [5] explored diffusion models and identified their ability to memorize samples under certain conditions. They acknowledge differential privacy as the gold standard defense strategy, but did not provide completed experiments to this end. Therefore, we evaluate the effectiveness of DP to this end, assessing memorization within our DP-trained models to investigate whether inadvertent reproduction of the training data can be eliminated.
To conduct this study, we adopt the evaluation methodology established by [5] to scrutinize the occurrence of memorization. Specifically, from each DP-trained diffusion model, we generate a vast number of samples (five times the size of the training set). Subsequently, for each generated image, we assess potential memorization compared to the original training samples using the adaptive distance metric introduced by [5],
$\ell\left(\hat{x},x;S_{\hat{x}}\right)=\frac{\ell_{2}(\hat{x},x)}{\alpha\cdot\mathbb{E}_{y\in S{\hat{x}}}\left[\ell_{2}(\hat{x},y)\right]}.$ (4)
Here, $S_{\hat{x}}$ denotes the set comprising the $n$ nearest elements from the training dataset to the example $\hat{x}$. The resulting distance metric yields a small value if the extracted image $x$ exhibits significantly closer proximity to the training image $\hat{x}$ compared to the $n$ closest neighbors of $\hat{x}$ within the training set. The idea is to find generated images that are unusually close to an original training image as indication of memorization. We set $\alpha=0.5$ and $n=50$ as in [5].
[5] did not define the specific threshold for Equation 4 for marking when a sample is considered memorized. Therefore, we consider the intuitive threshold to be less than 1, as this would indicate that the distance from the extracted image to the training image is less than half of the average distance to the closest $n$ neighbors. Upon conducting this assessment, we do not find any instances of memorized samples for all datasets under such definition, even at an elevated privacy parameter of $\epsilon=50$, with the closest distance values being $\sim$1.3. We show the histogram of scores for all samples on each dataset in Figure 3.
@@ 5.2 Addressing Memorization
Figure 3: Histogram of distance scores for all generated samples at $\epsilon=50$ to corresponding closest training image by Eq. 4 on each dataset. Note that the y-axis in in log scale, as there are very few samples with lower scores.
Because the threshold definition for memorization could vary, we also qualitatively show the samples with the lowest distances for all datasets at $\epsilon=50$ in Figure 4. Notably, the training versus the generated samples have discernible differences, in contrast to the nearly identical samples uncovered in [5] when training large diffusion models without DP. Also, given the nature of FL, the choice of diffusion model size will typically be small (for example, ours is $\sim$5.8M parameters), and therefore will be less likely to memorize compared to the larger DMs evaluated in [5]. As DP algorithms improve, we anticipate that even better final accuracy can be achieved while maintaining guaranteed privacy in the future with FedDiff.
@@ 5.3 Fourier Magnitude Filtering
In order to understand the impact of prioritizing data quality on performance, we conducted an initial experiment. For the CIFAR-10 dataset, we leverage a centralized pretrained classifier as an oracle to discern high and low quality samples. Specifically, we selectively retain samples for which the oracle accurately classifies and discarded those it misclassifies. This provides a way to filter out samples that are likely irrelevant or misleading for training a model. We then train the global model exclusively on the curated dataset of accurate samples and evaluate. This investigation yields a discernible improvement ranging from approximately 2% to 4% in final global model accuracy compared to training with all generated data, verifying an importance for data quality. Therefore, a critical question arises from this observation: how can we conduct sample filtration in the absence of an oracle?
To do so, we look to the Fourier domain for a potential source of information. As inspiration, we note that the use of the magnitude of local client images has been utilized in FL to assist in domain generalization across clients by providing low-level “style" information without the high-level semantics encoded in the phase [23]. In our case, we propose to leverage the Fourier magnitude information as a potential referenceable indicator to guide the sample filtering process. Furthermore, we are able to do so under very tight DP guarantees.
Specifically, on the client, we take the Fourier transform of the local samples and extract the magnitude information. For each client $c$, we gather the average sample magnitude with
$\bar{M_{c}}=\frac{1}{n_{c}}\sum_{i=1}^{n_{c}}|\psi(\mathbf{x}^{i})|,$ (5)
where $\psi$ is the 2D Fourier transform operation, $\mathbf{x}^{i}$ is a sample, and $n_{c}$ is the total number of samples in client $c$. $\bar{M}$ is bundled with the model and transmitted by the client to the relevant global party.
As in our standard global training procedure, samples are generated with the client-trained diffusion models to form a synthetic set. Prior to conducting global training, we calculate a sample score $s$ for the generated data $z$ from each diffusion model from the clients, $s_{\mathbf{z}^{i}_{c}}=\||\psi(\mathbf{z}^{i}_{c})|-\bar{M}_{c}\|_{2}$. We can then leverage this information to guide the removal of irrelevant samples, forming the final training set $\mathbf{G}$ by removing $\gamma$ percent of the generated data with the highest $s$ (larger magnitude difference). To continually ensure privacy guarantees, we apply DP in the FMF calculation. We do so by employing the DP bounded mean [19] from PyDP [^1] https://github.com/OpenMined/PyDP to calculate the average magnitude $\bar{M_{c}}$ at each client. This allows us to precisely manage any degree of privacy leakage for $\bar{M_{c}}$ and include it in the overall privacy budget.
@@ 6 Conclusion
In conclusion, our work addresses two valuable research questions in one-shot FL. Firstly, we investigate the potential of diffusion models for one-shot FL, and present the pioneering effort in this direction. In our investigation, we unveil the unique advantages that DMs offer, showcasing their potential to enhance the overall performance and tackle heterogeneity across diverse settings with our proposed approach, FedDiff. Secondly, we study privacy in SOTA one-shot FL and contribute a thorough investigation under provable privacy budgets, as well as address memorization concerns. Furthermore, to enhance performance under harsh DP conditions, we propose a novel and pragmatic solution, Fourier Magnitude Filtering, to boost the efficacy of generated data for global model training by eliminating low-quality samples. More discussions, including limitations and broader impact, are included in the Supplementary Material. We hope our work will inspire the community and foster further research in this direction to improve one-shot FL with generative models.

### v2

@@ 4.1 Data Heterogeneity
In Table 1, we analyze the performance of all methods under moderate ($Dir(0.1)$) to extreme ($Dir(0.001$) heterogeneity. Interestingly, FedDiff outperforms all other methods by a significant margin, from $\sim$5to up to $\sim$20in different scenarios. In the case of CIFAR-10, which is the most complex of the datasets, we find that FedDiff provides the most improvement. As discussed in our initial motivations (Sec. 3), we reason that the focus on sample quality and diversity that is provided by the DM objective enables much-improved performance. Intuitively, this becomes increasingly evident in more complex settings. To verify this observation, we conduct a comparative analysis of generated samples produced by FedCVAE and our FedDiff approach, illustrated in Figure 2. The discernible disparity is evident, with the samples generated by our method exhibiting significantly enhanced sharpness and overall quality.
@@ 4.2 Number of Clients
We also study the effect of the number of clients $C$ in Table 2. As we employ the same total number of samples in all experiments, the number of samples per client will increase with smaller $C$, and decrease with larger $C$. This allows us to observe the effect of increasing the distributed nature of the data across the client network.
One question arising from the adoption of generative models in FL settings pertains to their ability to maintain satisfactory performance when trained on a limited number of samples. Interestingly, when analyzing the results, we find that FedDiff is capable of handling a much smaller number of client training samples with little to no performance degradation. On the other hand, the discriminative model approaches quickly experience a collapse in performance when expanding to 20 clients. In the heterogeneous environment of federated learning, the local optimization of a discriminative model on a highly imbalanced and small dataset proves challenging. Rather than being an overwhelming burden, such a situation is handled well by FedDiff, as its sole focus is to capture the subsequently smaller distribution. Furthermore, we again find that FedDiff outperforms FedCVAE in all settings, further illustrating the potential for diffusion models in one-shot FL.
@@ 4.3 Resource Requirements
It is pertinent to emphasize that training diffusion models within the FedDiff framework is no more intricate than conventional methodologies and remains highly viable for FL. The computational complexity aligns with training a conventional CNN model with a modest number of parameters, and we employ the same number of local epochs as previous work with CNNs [38]. Importantly, the training process entails selecting random steps in the diffusion process at any given training iteration, eliminating the necessity for sequential steps during training. During inference, the generation process involves sequential denoising steps; however, this poses no issue for the clients, as generation occurs at the server in FedDiff. Therefore, FedDiff is an effective and practical approach for providing strong performance. Further discussions on communication and server-side operations are provided in the Supplementary Material.
@@ 5 RQ2: Privacy Considerations
Privacy holds paramount importance in one-shot FL. The trained client model may be repeatedly utilized or even exchanged in a model market context, and therefore, safeguarding the privacy of the model before it leaves the client is imperative. However, other SOTA works have not experimented with DP constraints, nor have they thoroughly explored this aspect, often leaving privacy discussions simply as a possibility for future work [38, 14]. In the subsequent sections, we meticulously investigate privacy from various perspectives and dive into our research question RQ2. Furthermore, we propose a novel approach in Sec. 5.3 to improve performance under strict privacy constraints.
@@ 5.1 Differential Privacy
Differential privacy is the widely accepted standard for ensuring privacy of a model, as it offers a provable guarantee of privacy [1, 8, 9, 7, 12]. Utilizing $(\epsilon,\delta)$ DP during model training guarantees comprehensive privacy protection, encompassing not only the model’s parameters and activations but also extending to all subsequent downstream operations. It is important to note that a model trained under $(\epsilon,\delta)$ DP safeguards the privacy of every training sample, regardless of its qualities or uniqueness [37]. Specifically, we train all approaches under ($\epsilon,\delta$) DP at the clients for various privacy levels of $\epsilon$ = 50, 25, and 10, with $\delta=10e^{-5}$, $C=10$, and $\alpha=0.01$. Lower $\epsilon$ values correspond to a tighter privacy budget, and the stated budget is for the entire training of each local model. We employ the Opacus [37] library for implementing DP. Further DP training details are provided in Supp. Material. We present the results in Table 4.
@@
As expected, all methods experience a drop in performance when trained under DP settings. Nonetheless, FedDiff still stands out, outperforming all other methods by a significant margin. Particularly for FashionMNIST, FedDiff experiences comparatively less accuracy drop under DP than FedCVAE. As articulated in our initial motivations outlined in Section 3, DP training introduces noise into the training process, exacerbating the complexity of optimization. In such scenarios, the simplicity of the training paradigm employed by diffusion models becomes notably advantageous. To demonstrate this even further, we conduct an experiment with an even tighter privacy budget of $\epsilon=1$ for FedDiff, achieving $65.53\pm 0.70$, $44.38\pm 3.35$, and $21.48\pm 1.53$ on FashionMNIST, PathMNIST, and CIFAR-10 respectively. As other methods already have very low accuracy at $\epsilon=10$, restricting any further would be counterproductive. On the other hand, despite facing even more stringent privacy constraints, FedDiff maintains a higher level of performance at $\epsilon$ = 1 than all other methods in Table 4 at $\epsilon$ = 50. Overall, we show that FedDiff is a strong approach even when DP is employed.
@@
Figure 3: Histogram of distance scores for all generated samples at $\epsilon=50$ to corresponding closest training image by Eq. 4 on each dataset. Note that the y-axis is in log scale, as there are very few samples with lower scores.
@@
5.2 Addressing Memorization
In federated learning, one of the most common attacks to violate the privacy of client models and reconstruct their local training data is to use gradient inversion [36, 10], where gradient information from the trained client model is exploited. However, existing literature lacks studies on performing gradient inversion on diffusion models, designing their attacks for discriminative architectures. Nonetheless, this omission is well-founded, as a more logical approach is to generate data directly via diffusion model sampling. Therefore, a crucial question arises: do DMs trained in a one-shot FL setting memorize training samples and regenerate them during inference? A recent study by [5] explored diffusion models in centralized settings and identified their potential to memorize samples under specific conditions. They acknowledged differential privacy as the gold standard defense strategy, but did not provide completed experiments. Therefore, we evaluate the effectiveness of DP to this end, assessing memorization within our DP-trained models to investigate whether inadvertent reproduction of the training data can be eliminated. To conduct this study, we adopt the evaluation methodology established by [5] to scrutinize the occurrence of memorization. Specifically, from each DP-trained diffusion model, we generate a vast number of samples (five times the size of the training set). Subsequently, for each generated image, we assess potential memorization compared to the original training samples using the adaptive distance metric introduced by [5],
$\ell\left(\hat{x},x;S_{\hat{x}}\right)=\frac{\ell_{2}(\hat{x},x)}{\alpha\cdot\mathbb{E}_{y\in S{\hat{x}}}\left[\ell_{2}(\hat{x},y)\right]}.\vspace{-2mm}$ (4)
Here, $S_{\hat{x}}$ denotes the set comprising the $n$ nearest elements from the training dataset to the example $\hat{x}$. The resulting distance metric yields a small value if the extracted image $x$ exhibits significantly closer proximity to the training image $\hat{x}$ compared to the $n$ closest neighbors of $\hat{x}$ within the training set. The idea is to find generated images that are unusually close to an original training image as an indication of memorization. We set $\alpha=0.5$ and $n=50$ as in [5]. [5] did not define the specific threshold for Equation 4 for marking when a sample is considered memorized. Therefore, we consider the intuitive threshold to be less than 1, as this would indicate that the distance from the extracted image to the training image is less than half of the average distance to the closest $n$ neighbors. Upon conducting this assessment, we do not find any instances of memorized samples for all datasets under such definition, even at an elevated privacy parameter of $\epsilon=50$, with the closest distance values being $\sim$1.3. We show the histogram of scores for all samples on each dataset in Figure 3. Because the threshold definition for memorization could vary, we also qualitatively show the samples with the lowest distances for all datasets at $\epsilon=50$ in Figure 4. Notably, the training versus the generated samples have discernible differences, in contrast to the nearly identical samples uncovered in [5] when training large DMs without DP. Also, given the nature of FL, the choice of diffusion model size will typically be small (for example, ours is $\sim$5.8M parameters), and therefore will be less likely to memorize compared to the larger DMs evaluated in [5].
@@ 5.3 Fourier Magnitude Filtering
To understand the impact of prioritizing data quality on performance, we conduct an initial experiment. For CIFAR-10, we leverage a centralized pretrained classifier as an oracle to discern high and low quality samples. We selectively retain samples for which the oracle accurately classifies and discard those it misclassifies. This provides a way to filter out samples that are likely irrelevant or misleading for training a model. We then train the global model exclusively on the curated dataset of accurate samples and evaluate. This investigation yields a discernible improvement ranging from approximately 2% to 4% in final global model accuracy compared to training with all generated data, verifying the importance of data quality. Therefore, a critical question arises from this observation: how can we conduct sample filtration in the absence of an oracle?
To do so, we look to the Fourier domain for a potential source of information. We note that the magnitude of local client images has been utilized in FL to assist in domain generalization across clients by providing low-level “style” information without the high-level semantics encoded in the phase [24]. In our case, we propose to leverage the Fourier magnitude information as a potential referenceable indicator to guide the sample filtering process.
On the client, we take the Fourier transform of the local samples and extract the magnitude information. For each client $c$, we gather the average sample magnitude with
$\bar{M_{c}}=\frac{1}{n_{c}}\sum_{i=1}^{n_{c}}|\psi(\mathbf{x}^{i})|,\vspace{-2mm}$ (5)
where $\psi$ is the 2D Fourier transform operation, $\mathbf{x}^{i}$ is a sample, and $n_{c}$ is the total number of samples in client $c$. $\bar{M}$ is bundled with the model and transmitted by the client to the relevant global party. As in our standard global training procedure, samples are generated with the client-trained diffusion models to form a synthetic set. Prior to conducting global training, we calculate a sample score $s$ for the generated data $z$ from each diffusion model from the clients, $s_{\mathbf{z}^{i}_{c}}=\||\psi(\mathbf{z}^{i}_{c})|-\bar{M}_{c}\|_{2}$. We can then leverage this information to guide the removal of irrelevant samples, forming the final training set $\mathbf{G}$ by removing $\gamma$ percent of the generated data with the highest $s$ (larger magnitude difference). To continually ensure privacy guarantees, we apply DP in the FMF calculation. We do so by employing the DP bounded mean [20] from PyDP [27] to calculate the average magnitude $\bar{M_{c}}$ at each client. This allows us to precisely manage any degree of privacy leakage for $\bar{M_{c}}$ and include it in the overall privacy budget.
@@ 6 Conclusion
Our work addresses two valuable research questions in one-shot FL. Firstly, we investigate the potential of DMs for one-shot FL and present the pioneering effort in this direction. We unveil the unique advantages that DMs offer, showcasing their potential to enhance performance and tackle heterogeneity across diverse settings with our proposed approach, FedDiff. Secondly, we study privacy in SOTA one-shot FL and contribute a thorough investigation under provable privacy budgets, as well as address memorization concerns. Furthermore, to enhance performance under harsh DP conditions, we propose a novel and pragmatic solution, Fourier Magnitude Filtering, to boost the efficacy of generated data for global model training by eliminating low-quality samples. More discussions, including limitations and broader impact, and additional experiments are included in the Supp. Material. We hope our work will inspire the community and foster further research in this direction to improve one-shot FL with generative models.
Acknowledgment. This work is partially supported by the NSF/Intel Partnership on MLWiNS under Grant No. 2003198.
Supplementary Material
Appendix A Additional Training Details
The FashionMNIST dataset is an alternative to the original MNIST dataset, providing a more challenging task by replacing the handwritten digits with grayscale images of various fashion items. The dataset consists of 60,000 training images and 10,000 test images. The PathMNIST dataset is a medical dataset of colon pathology images in RGB, with a training set of 89,996 images and a test set containing 7,180 images with 9 classes. The CIFAR-10 dataset consists of 60,000 color images equally distributed into ten different classes. The dataset is composed of a training set containing 50,000 images and a test set comprising of 10,000 images. CIFAR-10 is natively sized at 32$\times$32 pixels. We upsample FashionMNIST and PathMNIST from 28$\times$28 to 32$\times$32. A visualization of the dataset partitioning across clients is shown in Figure 6.
We train with a batch size of 128 for all methods and use the AdamW optimizer. For local (and global training were applicable), we searched learning rates from [$3e^{-3}$, $1e^{-3}$, $3e^{-4}$, $1e^{-4}$] for each method using the CIFAR-10 dataset to find the optimal settings. We employed a ResNet16 architecture for the global model of all methods to ensure a fair comparison. For DP experiments, we set the max gradient norm clipping threshold to 1.0 for all experiments and methods. In accordance with the recommendations of the Opacus [37] library, we employ their Poisson batch sampling to ensure privacy guarantees.
As mentioned in Section 3.1 of the main paper, our DM is a basic U-Net structure with residual blocks [15, 29] and class-conditioning. Specifically, our U-Net has three downsampling stages (1/8 total downsampling) and three upsampling stages, each with residual convolutional blocks. A learnable embedding for time step and class conditioning is concatenated as additional input channels. For FedDiffS, we halve the number of channels per block to reduce model size. For sampling at the server, we perform 1000 iterations as in [15] to generate each batch. The total number of generated samples is set equal to the size of the original dataset. Code available at https://github.com/mmendiet/FedDiff.
(a) $\alpha=0.1$
(b) $\alpha=0.001$
Figure 6: $Dir(\alpha)$ data partitioning for 10 clients on CIFAR-10. We show moderate ($\alpha=0.1$) to severe ($\alpha=0.01$) data heterogeneity levels. Data heterogeneity poses a significant challenge for many one-shot FL methods, as reconciling various models trained on widely different distributions is non-trivial. Our FedDiff approach rather trains diffusion models on the simple client distributions, which can then generate useful synthetic data for training global models.
(a) FashionMNIST
(b) PathMNSIT
(c) CIFAR-10
Figure 7: Ablation study of $\gamma$ in FMF under the $\epsilon$ = 10 setting. The accuracy of FedDiff is in green and FedDiff+FMF for various $\gamma$ in blue. Generally, data filtering within the range of 1% to 10% produces positive outcomes, resulting in improved performance, with approximately 5% serving as an effective default choice. We plot the mean across three runs with different seeds for each setting.
Appendix B Communication and Server-side Operations
All methods primarily involve transmitting the model weights to the server, and this is done a single time. Additionally, specific information is sent. FedAvg and DENSE transmit the number of samples and label space, while FedCVAE and FedDiff send the number of samples per label. Thus, communication cost is mainly determined by the model size and number of clients. To ensure fairness between generative and discriminative methods, we select models with similar parameters and FLOPs, as shown in Table 3, maintaining comparable communication and computation costs. We describe details of our DM architectures in Section A. As we also show in our experiments with FedDiffS in Table 3 of the main paper, we can adjust the size of the model to meet communication or compute needs and still provide exceptional performance.
Once the models are on the server, different operations are required for each method. DENSE, FedCVAE, and FedDiff require server-side generation and training, producing a dataset the same size as the original (e.g., 50k for CIFAR-10) and training a global model for 200 epochs. Server-side operation times on a single A5000 GPU for CIFAR-10 with 10 clients are 3.36, 1.76, and 2.17 hours for DENSE, FedCVAE, and FedDiff, respectively. DENSE takes the longest due to nested training of a GAN and the global model. Nonetheless, server computation is not a primary concern in FL, as it is not constrained by the same resource limitations as client devices.
Appendix C Sampling Steps for Generation Ablation
Diffusion models generate images through consecutive denoising steps, making the number of iterations per image a design choice when generating synthetic datasets at the server. Table 5 provides insights into the impact of adjusting this parameter. Generally, increasing the number of steps enhances data quality and improves the final global model performance, particularly for the more difficult datasets. However, if generation time on the server is a concern, this parameter can be reduced, or increased when prioritizing data quality and model accuracy.
Table 5: Ablation on number of diffusion steps ($S$) used per image when generating the global synthetic dataset. Final global model accuracy is reported using our FedDiff approach under the default setting of $\alpha=0.01$ and $C=10$. The 1000-step setting is employed across all other experiments in the paper, as it is the standard practice in DDPM [15].
Dataset $S=100$ $S=500$ $S=1000$ $S=2000$
FashionMNIST 86.49$\pm$0.23 86.63$\pm$0.38 86.81$\pm$0.54 86.75$\pm$0.17
PathMNIST 69.85$\pm$1.15 70.13$\pm$1.61 70.61$\pm$1.37 71.44$\pm$1.83
CIFAR-10 55.89$\pm$1.93 56.13$\pm$1.74 56.57$\pm$2.42 57.74$\pm$2.86
Appendix D FMF $\gamma$ Ablation
In Figure 7, we present the outcomes obtained using FedDiff+FMF under $\epsilon=10$ across a range of $\gamma$ values, encompassing data filtering percentages spanning from 1% to 12%. Our findings indicate that, in general, data filtering within the 1% to 10% range yields favorable results and leads to performance enhancements, with around 5% being a great default. Interestingly, the degree of improvement provided by FMF becomes more pronounced and consistent as the dataset becomes more challenging. This phenomenon aligns with the anticipated trends, as more intricate datasets inherently pose a greater challenge, making it less likely for the generators to consistently produce high-quality samples. Consequently, the need for data filtering becomes more pronounced in such scenarios to enhance sample quality. This trend is also favorable since it addresses the specific need for improvement, especially in cases where performance is suboptimal and the challenges are more pronounced.
Appendix E Discussions, Limitations and Broader Impact
Model Heterogeneity. In real FL systems, model heterogeneity may often occur [38, 14]. For instance, some clients may have architecture variations in their models or have smaller or larger models depending on their computing capabilities. Therefore, clients may have different architectures of similar generation capability, or even differing capabilities depending on the requirements of each client. Our approach allows for flexibility to accommodate such system diversity across clients. In FedDiff, we generate data from the client models and employ that synthetic data for global training, and therefore can leverage varying models without the worry of reconciling the weights themselves.
Limitations and Broader Impact. One downside of our method is that the generated data, particularly under DP constraints, still lacks in quality and effectiveness for global model training versus using true data. For instance, with DP on CIFAR-10 as shown in Figure 4 in the main paper, the data loses a substantial amount of structure. An interesting direction for future work would be to study how to further improve the quality of the generated data and its usefulness for global model training while maintaining privacy. For instance, as differential privacy algorithms improve, FedDiff and the generated data quality will likewise benefit. Additionally, potentially leveraging prior information could allow the models to focus on task-relevant features, excluding irrelevant ones. By training a separate model to identify important features and then learning to generate images based only on these features with a DP secure model, we could potentially simplify the information needed for the generation process. This would likely lead to quicker convergence, allowing for better learning in DP settings with the same privacy budget and ultimately improving the privacy-utility trade-off of the generated samples.
Looking at the broader impact of our work, FL depends on the diversity of data contributed by different participants. If biases exist in the local datasets, they can be propagated and amplified during the model training process. This could lead to unintended algorithmic biases and discrimination in the resulting models. Ensuring diversity and fairness in the data used for FL is an important research direction to mitigate this risk and promote equitable outcomes [2], particularly in the highly data heterogeneous environments explored in this work. Furthermore, as we have discussed throughout our paper, the privacy of client data is important in FL. To mitigate risks in this regard, we take many precautions to preserve privacy of the clients participated in the FL process though the use of DP, and operating within the one-shot setting to reduce the chance of eavesdropping.